# Adaptive Label Correction for Robust Medical Image Segmentation with Noisy Labels

### v1

@@ III-D Ablation Studies
To further evaluate the label refinement method, we compared the quality of directly generated pseudo-labels with refined pseudo-labels for the top-$k$ samples, evaluated every 5,000 iterations. The results, visualized in Fig. 3, demonstrate that although there are occasional fluctuations where generated pseudo-labels outperform refined ones, label refinement consistently enhances overall performance. On the LA and NIH datasets, refined labels generally outperform directly generated pseudo-labels. This improvement is attributed to label refinement’s ability to effectively capture perturbation-induced relationships, yielding more accurate pseudo-labels.
III-E Parameter Sensitivity Analysis
To evaluate the impact of different values of $k$, experiments were conducted on the LA and NIH datasets, as shown in Fig. 5. The parameter $k$ controls the proportion of low-quality labeled data in a mini-batch used for gradient updates. A smaller $k$ reduces the inclusion of refined labels, offering limited benefits to the model, while a larger $k$ increases the likelihood of incorporating lower-quality refined labels, potentially leading to performance degradation. Optimal performance is achieved at $k=0.5$, which balances label quality and data utilization effectively.
Additionally, we analyzed the influence of $\alpha$ and $\beta$ on the NIH pancreas dataset, as shown in Fig. 6. These parameters modulate the contributions of $\mathcal{L}_{ls}$ and $\mathcal{L}_{n}$, respectively. Larger values of $\alpha$ and $\beta$ can destabilize training by overemphasizing high-quality labels, while smaller values may hinder performance improvements. Optimal results are achieved at $\alpha=3$ and $\beta=2$, ensuring effective supervision and stable training.
IV Conclusion

### v3

@@ IV-D Ablation Studies
IV-E Parameter Sensitivity Analysis
To assess the robustness and generalizability of our method, we conducted extensive sensitivity analysis on key hyperparameters involved in our framework. Specifically, we focus on two aspects: the sample selection ratio $k$, and the loss weighting parameters $\alpha$ and $\beta$.
Impact of $k$ on Training Dynamics. The parameter $k$ controls the proportion of low-quality (LQ) labeled samples used in each mini-batch. As shown in Fig. 5, both overly small and large $k$ values degrade performance. Small $k$ leads to underutilization of useful LQ data, while large $k$ increases the risk of incorporating unreliable refined labels, introducing noise during optimization. The best performance is achieved at $k=0.5$, which effectively balances label quality and data diversity. This demonstrates the importance of selective sample inclusion and confirms the robustness of our label selection strategy.
Influence of $\alpha$ and $\beta$. We also analyzed the effects of $\alpha$ and $\beta$, which weight the label selection loss $\mathcal{L}{ls}$ and noisy label loss $\mathcal{L}{n}$, respectively. A grid search over $\alpha,\beta\in[1,5]$ on the NIH dataset under 10% HQ / 90% LQ conditions yielded the results shown in Fig. 6. The surface plot reveals nonlinear sensitivity, where both overly low and high values reduce Dice performance, confirming the need for balanced loss weighting. The optimal region centers around $\alpha=3$, $\beta=2$, with stable performance in nearby ranges (e.g., $\alpha\in[2.5,3.5]$). In contrast, high values (>4>4) lead to overfitting due to excessive reliance on HQ labels, while low values (<1.5<1.5) weaken the benefits of adaptive supervision. Overall, these results validate the effectiveness and stability of our loss design. Future work may explore adaptive or learnable weighting schemes to further enhance flexibility and performance across varying noise conditions.
V Conclusion